# A deep neural network approach to solve the Dirac equation

### v1

@@ I Introduction
The pioneer work solved spin systems by using the Boltzmann machine [2]. The following work extended this method with a higher accuracy with introducing pair-wise ansatz [3] and deterministic time evolution [4]. The extension to the excited-state calculation with two different ansatz were proposed in Refs. [5, 6]. Other approaches to solve spin systems applied Jastrow-Slater wave function ansatz [7] and hidden-fermion one [8]. Reference [9] proposed a generalization model for spins, which is to take perceptrons as the transfer matrix of a spin-glass Hamiltonian.
For continuous systems, the design of the machine learning algorithm should be based on the (anti)symmetrization of the system. Several models of bosonic systems were solved with a fully connected deep neural network (DNN) in Ref. [10]. Reference [11] employed the Deep Sets neural-network architecture to solve periodic systems. For fermionic systems, from electrons to nucleons, the Jastrow-Slater ansatz was a successful application to consider the antisymmetrization [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Recent progress on this direction is summarized in Ref. [26]. Another application was based on the Jordan-Wigner mapping or similar mapping methods from the fermionic system into a spin system and using the Boltzmann machine [27]. Reference [28] proposed a DNN model to solve the ground state and excited states in the coordinate space, which included symmetrization for bosons or antisymmetrization for fermions. Soon this method was extended with spin and isospin degrees of freedom by applying a non-fully connected DNN structure [29]. However, all these works focused on the non-relativistic case. The relativistic formulation using the Dirac equation is sometimes crucial to describe electronic structure of atoms [30], molecules [31], and nuclear systems [32]. However, the unsupervised machine learning has not been applied to the Dirac equation yet; one of the difficulties of the Dirac equation using a variational method is the so-called the variational collapse. Due to the existence of the Dirac sea, the ground state energy is not the lowest eigenvalue of a Dirac Hamiltonian. Therefore, minimizing the energy directly causes the loss function falling into the Dirac sea, which is the variational collapse.
@@
This paper is organized as follows: In Sec. II, we discuss the Dirac equation with spherical symmetry, and introduce the inverse Hamiltonian method to solve the ground state, where the excited states are solved by both the inverse Hamiltonian and the orthonormal method. In Sec. III, we introduce our DNN model. The results of Coulomb and Wood-Saxon potential calculations are shown in Secs. IV and V, respectively. Section VI is the summary of our work.
@@ II.1 The Dirac equation with spherical symmetry
where $F_{n\kappa}\left(r\right)$ and $G_{n\kappa}\left(r\right)$ denote the large and small components of a radial wave function, respectively. The subscripts $n$ and $j_{z}$, respectively, denote the principal quantum number the $z$-projection of the total angular momentum $j$. For a given $j$, $\kappa=j+1/2$ if the orbital angular momentum $l$ satisfies $l=j+1/2$, and $\kappa=-j-1/2$ if $l$ satisfies $l=j-1/2$. The detailed forms of the tensor spherical harmonics are given by
@@
where $H^{\prime}_{\scriptstyle{\text{{{{D}}}}}}$ and $\varphi_{n\kappa}$, respectively, read
@@ II.2 The inverse Hamiltonian method
The ground state is $\kappa=-1$ and $n=1$; thus, we choose an arbitrary value between the ground-state energy $\varepsilon_{1,-1}$ and the highest eigenvalue of negative energy states $\left(\varepsilon_{1,-1}^{*}-2mc^{2}\right)$ for $\varepsilon^{\prime}_{1,-1}$, that is, $\left(\varepsilon_{1,-1}^{*}-2mc^{2}\right)<\varepsilon^{\prime}_{1,-1}<\varepsilon_{1,-1}$ to obtain the ground-state energy. The schematic figure of the spectrums of $H^{\prime}_{\scriptstyle{\text{{{{D}}}}}}$ and $H_{\scriptstyle{\text{{{{inv}}}}}}^{1,-1}$ is shown in Fig. 1. The lowest eigenstate of $H_{\scriptstyle{\text{{{{inv}}}}}}^{1,-1}$ corresponds the ground state.
Another benefit of applying the inverse Hamiltonian method is that both the ground and low-lying excited states can be calculated by varying $\varepsilon^{\prime}_{n\kappa}$. To calculate the state of quantum number $n+1$ and $\kappa$, we set $\varepsilon^{\prime}_{\left(n+1\right)\kappa}$ satisfying between
$\varepsilon_{n\kappa}<\varepsilon^{\prime}_{\left(n+1\right)\kappa}<\varepsilon_{\left(n+1\right)\kappa}.$ (14)
@@
$\varepsilon_{\left(n+1\right)\kappa}=\varepsilon^{\prime}_{\left(n+1\right)\kappa}-\left(\min\frac{\left\langle{\varphi}\middle|{H_{\scriptstyle{\text{{{{inv}}}}}}^{\left(n+1\right)\kappa}}\middle|{\varphi}\right\rangle}{\left\langle{\varphi}\middle|{\varphi}\right\rangle}\right)^{-1}.$ (15)
Practically, $\varepsilon^{\prime}$ is chosen slightly smaller than zero. If the trial with another parameter $\varepsilon^{\prime\prime}$ which is much smaller than $\varepsilon^{\prime}$ returns the same eigenvalue, we can safely assume the obtained state is the ground state. If the test with $\varepsilon^{\prime\prime}$ returns a lower eigenvalue, then we repeat the process until we obtain the ground state with two different trial parameters with a large gap. In the case with spherical symmetry, a more direct strategy is to count the number of nodes in the big the component of wave functions, which is equal to $n$. Then, in a recursive approach, the state of quantum number $n+1$ and $\kappa$ can be obtained by setting $\varepsilon^{\prime}_{\left(n+1\right)\kappa}$ slightly greater than the $n$-th eigenvalue $\varepsilon_{n\kappa}$.
Figure 1: (a) Spectrum of eigenstates of the Dirac Hamiltonian $H^{\prime}_{\scriptstyle{\text{{{{D}}}}}r}$. The half-filled blue circles denote the ground state and low-lying excited states, while the blue line denotes positive continuum states. The half-filled red circles and red line denote the bound states and continuum states in the Dirac sea, respectively. (b) Spectrum of eigenstates of the inverse Hamiltonian $H_{\scriptstyle{\text{{{{inv}}}}}}^{1,-1}$ [33] for the ground state calculation. The continuum states are reversed in the middle between bound states. The order among the bound states in both the Fermi sea and the Dirac sea does not change, respectively. The ground state is the lowest state in the spectrum.
@@ II.3 The orthonormal method
$\varepsilon_{\left(n+1\right)\kappa}=\varepsilon^{\prime}_{n_{\scriptscriptstyle{\text{{{{min}}}}}}\kappa}-\left(\min\frac{\left\langle{\varphi^{\prime}}\middle|{H_{\scriptstyle{\text{{{{inv}}}}}}^{n_{\scriptscriptstyle{\text{{{{min}}}}}}\kappa}}\middle|{\varphi^{\prime}}\right\rangle}{\left\langle{\varphi^{\prime}}\middle|{\varphi^{\prime}}\right\rangle}\right)^{-1}.$ (18)
@@ III.1 Discretization of the Dirac equation
$V\pm S\simeq\begin{pmatrix}\tilde{V}_{1}\pm\tilde{S}_{1}&0&0&\cdots&0&0&0\\ 0&\tilde{V}_{2}\pm\tilde{S}_{2}&0&\cdots&0&0&0\\ 0&0&\tilde{V}_{3}\pm\tilde{S}_{3}&\cdots&0&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&\tilde{V}_{M-3}\pm\tilde{S}_{M-3}&0&0\\ 0&0&0&\cdots&0&\tilde{V}_{M-2}\pm\tilde{S}_{M-2}&0\\ 0&0&0&\cdots&0&0&\tilde{V}_{M-1}\pm\tilde{S}_{M-1}\end{pmatrix},$ (21)
where $\tilde{V}_{i}=V\left(r_{i}\right)$ and $\tilde{S}_{i}=S\left(r_{i}\right)$. The form of the differential operator depends on the distribution of mesh points, which will be discussed in Sec. IV.2 and Sec. V.2. Then, the Dirac Hamiltonian $H^{\prime}_{\scriptstyle{\text{{{{D}}}}}r}$ is expressed by a ($22M-$ 2)-dimensional matrix, which is used to construct the inverse Hamiltonian $H_{\scriptstyle{\text{{{{inv}}}}}}^{n\kappa}$ numerically.
@@ III.2 Design of the neural network
is used as the activation function. The parameters of the DNN are updated by the Adam optimizer [36].
@@ IV.1 The exact results
We apply our DNN method for the hydrogen atom. The scalar and vector potentials, respectively, read
@@ IV.2 Discrete representations with the log mesh
The Coulomb potential changes rapidly in the central region (small $r$). In order to calculate the central region precisely, we define $x$ by
@@
and $x$ is discretized uniformly with the size $\Delta x$ and $x_{0}$ as the small boundary point. Note that the term $e^{x_{0}}$ is introduced to impose the boundary condition $f_{n\kappa}\left({r=0}\right)=0$ properly, as the following Eq. (30) requires $f_{n\kappa}\left(r\left(x_{0}\right)\right)=0$. The first-order differential operator for the log mesh reads
@@ IV.3 DNN results
The wave functions of the ground and first five excited states for $\kappa=-1$ are shown in Figs. 3 and 4. In general, both methods can reproduce the exact wave functions, except the peak of the $n=6$ state at large $r$, since the $100\,\mathrm{a.u.}$ box size is too small for the $n=6$ state. The relative errors between DNN wave functions and exact results are shown in Fig. 5. Except around the boundary, the DNN reproduces the exact wave functions. The precisions of wave functions are almost the same level for both methods. Note that due to the existence of nodes, the relative error diverges at the nodes.
The relative errors of the loss function to the exact energy as a function of epochs for both methods are shown in Fig. 6. As shown in Table 1, the number of epochs to converge for the orthonormal method is one order of magnitude less than the inverse Hamiltonian method, while the time per epoch of the orthonormal method is about twice longer than the inverse Hamiltonian method. The calculation cost of the inverse Hamiltonian method does not depend on $n$; in contrast, larger $n$ requires longer calculation due to the orthonormalization.
@@ V.2 Discrete representations with uniform mesh
The uniform mesh is adopted to calculate for the Woods-Saxon potential, where $r$ is uniformly distributed with mesh size $\Delta r$. The derivative matrix of the first-order differential operator $\frac{\partial}{\partial r}$ reads
@@ V.3 DNN results
In this section, we use a $20\,\mathrm{fm}$ box with $M=2000$ meshes for the DNN calculation. The general pesudospectral method [39, 40] is adopted to calculate benchmark wave functions and energies with the mapping parameter $L=10$ and the number of mesh points $600$.
@@
The $F$- and $G$- components for $\kappa=-1$ of ${}^{208}\mathrm{Pb}$ are shown in Figs. 9 and 10, respectively. Both methods show good agreements with the benchmark for the $F$-component. According to Eq. (17), the accuracy of the orthonormal method depends on the accuracy of lower-state wave functions, which means it also depends on the the inverse Hamiltonian method which provides the lowest state data.
@@
Table 2: Energies of ${}^{16}\mathrm{O}$ below the Fermi energy calculated by the Woods-Saxon potential. The benchmark is obtained by the general pesudospectral method.
@@
$\varepsilon^{\prime}_{1\kappa}$ ($\mathrm{MeV}$) $-45.0$ $-28.0$ $-20.0$
@@
Figure 8: Spectrum of neutrons calculated by the Woods-Saxon potential of ${}^{208}\mathrm{Pb}$. The benchmark is given by the general pesudospectral method and the inverse Hamiltonian method is used for the DNN.
@@ VI Summary
To verify the feasibility of our method, we applied our method to calculate a hydrogen atom and the Woods-Saxon potentials of ${}^{16}\mathrm{O}$ and ${}^{208}\mathrm{Pb}$. We found that both methods show good agreements with the benchmark with about the same accuracy (up to $0.15\,\%$ difference) for the both systems.
We believe one promising extension of our work is to solve the relativistic many-body systems. In the case of the two-body systems, the two-body Dirac equation [41] describes two particles with interaction, where our DNN method may help in finding the solution especially for the excited states.

### v2

@@ I Introduction
The pioneer work solved spin systems by using the Boltzmann machine [2]. Subsequent research enhanced this approach, achieving higher accuracy through the introduction of a pair-wise ansatz [3] and deterministic time evolution [4]. The extension to the excited-state calculation with two different ansatz were proposed in Refs. [5, 6]. Other approaches to solve spin systems included to apply Jastrow-Slater wave function ansatz [7] and hidden-fermion one [8]. Reference [9] proposed a generalization model for spins, which was to take perceptrons as the transfer matrix of a spin-glass Hamiltonian.
For continuous systems, the design of the machine learning algorithm should consider the (anti)symmetrization of the system. Several models of bosonic systems were solved with a fully connected deep neural network (DNN) in Ref. [10]. Reference [11] employed the Deep Sets neural-network architecture to solve periodic systems. For fermionic systems, from electrons to nucleons, the Jastrow-Slater ansatz was a successful application to consider the antisymmetrization [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Recent progress on this direction is summarized in Ref. [26]. Another application was based on the Jordan-Wigner mapping or similar mapping methods from the fermionic system into a spin system to use the Boltzmann machine [27]. Reference [28] proposed a DNN model to solve the ground state and excited states in the coordinate space, which included symmetrization for bosons or antisymmetrization for fermions. Soon this method was extended with spin and isospin degrees of freedom by applying a non-fully connected DNN structure [29]. However, all these works focused on the non-relativistic case except for Ref. [30], which employed physics-informed neural networks to solve the time-dependent Dirac equation. The relativistic formulation using the Dirac equation is sometimes crucial to describe electronic structure of atoms [31], molecules [32], and nuclear systems [33]. Notably, the unsupervised machine learning has not been applied to the Dirac equation yet; a fundamental challenge in variational treatments of the Dirac equation is variational collapse A fundamental challenge in variational treatments of the Dirac equation is variational collapse. Because the Dirac sea exists, the ground-state energy does not correspond to the lowest eigenvalue of the Dirac Hamiltonian. Direct energy minimization therefore causes the loss function to plunge into the Dirac sea—a phenomenon known as variational collapse.
@@
This paper is organized as follows: In Sec. II, we discuss the Dirac equation with spherical symmetry, and introduce the inverse Hamiltonian method to solve the ground state, while the excited states are solved by both the inverse Hamiltonian and the orthonormal method. In Sec. III, we introduce our DNN model. The results of Coulomb and Wood-Saxon potential are shown in Secs. IV and V, respectively. Section VI is the summary of our work.
@@ II.1 The Dirac equation with spherical symmetry
where $F_{n\kappa}\left(r\right)$ and $G_{n\kappa}\left(r\right)$ denote the large and small components of a radial wave function, respectively. The subscripts $n$ and $j_{z}$, respectively, denote the principal quantum number and the $z$-projection of the total angular momentum $j$. For a given $j$, $\kappa=j+1/2$ if the orbital angular momentum number $l$ satisfies $l=j+1/2$, and $\kappa=-j-1/2$ if $l$ satisfies $l=j-1/2$. The detailed forms of the tensor spherical harmonics are given by
@@
where $H^{\prime}_{\scriptstyle{\text{{{{D}}}}}r}$ and $\varphi_{n\kappa}$, respectively, read
@@ II.2 The inverse Hamiltonian method
The ground state is $\kappa=-1$ and $n=1$; thus, we choose an arbitrary value between the ground-state energy $\varepsilon_{1,-1}$ and the highest eigenvalue of negative energy states $\left(\varepsilon_{1,-1}-2mc^{2}\right)$ for $\varepsilon^{\prime}_{1,-1}$, that is, $\left(\varepsilon_{1,-1}-2mc^{2}\right)<\varepsilon^{\prime}_{1,-1}<\varepsilon_{1,-1}$ to obtain the ground-state energy. The schematic figure of the spectrums of $H^{\prime}_{\scriptstyle{\text{{{{D}}}}}r}$ and $H_{\scriptstyle{\text{{{{inv}}}}}}^{1,-1}$ is shown in Fig. 1. The lowest eigenstate of $H_{\scriptstyle{\text{{{{inv}}}}}}^{1,-1}$ corresponds the ground state.
Another benefit of applying the inverse Hamiltonian method is that both the ground and low-lying excited states can be calculated by varying $\varepsilon^{\prime}_{n\kappa}$. To calculate the state of quantum number $n$ and $\kappa$, we set $\varepsilon^{\prime}_{n\kappa}$ satisfying between
$\varepsilon_{\left(n-1\right)\kappa}<\varepsilon^{\prime}_{n\kappa}<\varepsilon_{n\kappa}.$ (14)
@@
$\varepsilon_{n\kappa}=\varepsilon^{\prime}_{n\kappa}-\left(\min\frac{\left\langle{\varphi}\middle|{H_{\scriptstyle{\text{{{{inv}}}}}}^{n\kappa}}\middle|{\varphi}\right\rangle}{\left\langle{\varphi}\middle|{\varphi}\right\rangle}\right)^{-1}.$ (15)
Practically, $\varepsilon^{\prime}$ is chosen slightly smaller than zero. If the trial with another parameter $\varepsilon^{\prime\prime}$ which is much smaller than $\varepsilon^{\prime}$ returns the same eigenvalue, we can safely assume the obtained state is the ground state. If the test with $\varepsilon^{\prime\prime}$ returns a lower eigenvalue, we repeat the process until the ground state is obtained for two distinct trial parameters with a significant energy gap. In the case with spherical symmetry, a more direct strategy is to count the number of nodes in the large component of wave functions, which is equal to $n-1-l$. Then, in a recursive approach, the state of quantum number $n$ and $\kappa$ can be obtained by setting $\varepsilon^{\prime}_{n\kappa}$ slightly greater than $\varepsilon_{\left(n-1\right)\kappa}$. The eigenvalue $\varepsilon_{\left(n-1\right)\kappa}$ and $\varepsilon_{n\kappa}$ can be splited as the highest and lowest eigenvalues in the spectrum of the inverse Hamiltonian. Therefore, the inverse Hamiltonian method is a robust method even for highly dense spectra. Note that the orthonormal method, which will be discussed in the next subsection, has proven in Ref. [28] that it enables us to calculate even the degenerated spectrum, i.e., extremely dense spectra.
Figure 1: (a) Spectrum of eigenstates of the Dirac Hamiltonian $H^{\prime}_{\scriptstyle{\text{{{{D}}}}}r}$. The half-filled blue circles denote the ground state and low-lying excited states, while the blue line denotes positive continuum states. The half-filled red circles and red line denote the bound states and continuum states in the Dirac sea, respectively. (b) Spectrum of eigenstates of the inverse Hamiltonian $H_{\scriptstyle{\text{{{{inv}}}}}}^{1,-1}$ [34] for the ground-state calculation. The continuum states are reversed in the middle between bound states. The order among the bound states in both the Fermi sea and the Dirac sea does not change, respectively. The ground state is the lowest state in the spectrum.
@@ II.3 The orthonormal method
$\varepsilon_{\left(n+1\right)\kappa}=\varepsilon^{\prime}_{1\kappa}-\left(\min\frac{\left\langle{\varphi^{\prime}}\middle|{H_{\scriptstyle{\text{{{{inv}}}}}}^{1\kappa}}\middle|{\varphi^{\prime}}\right\rangle}{\left\langle{\varphi^{\prime}}\middle|{\varphi^{\prime}}\right\rangle}\right)^{-1}.$ (18)
@@ III.1 Discretization of the Dirac equation
$V\pm S\simeq\begin{pmatrix}\tilde{V}_{1}\pm\tilde{S}_{1}&0&0&\cdots&0&0&0\\ 0&\tilde{V}_{2}\pm\tilde{S}_{2}&0&\cdots&0&0&0\\ 0&0&\tilde{V}_{3}\pm\tilde{S}_{3}&\cdots&0&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&\tilde{V}_{M-3}\pm\tilde{S}_{M-3}&0&0\\ 0&0&0&\cdots&0&\tilde{V}_{M-2}\pm\tilde{S}_{M-2}&0\\ 0&0&0&\cdots&0&0&\tilde{V}_{M-1}\pm\tilde{S}_{M-1}\end{pmatrix}$ (21)
with $\tilde{V}_{i}=V\left(r_{i}\right)$ and $\tilde{S}_{i}=S\left(r_{i}\right)$. The form of the differential operator depends on the distribution of mesh points, which will be discussed in Sec. IV.2 and Sec. V.2. Then, the Dirac Hamiltonian $H^{\prime}_{\scriptstyle{\text{{{{D}}}}}r}$ is expressed by a ($22M-$ 2)-dimensional matrix, which is used to construct the inverse Hamiltonian $H_{\scriptstyle{\text{{{{inv}}}}}}^{n\kappa}$ numerically.
@@ III.2 Design of the neural network
is used as the activation function. The parameters of the DNN are updated by the Adam optimizer [37] with a fixed learning rate of 0.001. Regularization techniques are not applied, as no overfitting is observed during the trainings.
@@ IV.1 The exact results
We apply our DNN method for the hydrogen atom. The vector and scalar potentials, respectively, read
@@ IV.2 Discrete representations with the log mesh
The potential changes rapidly in the central region (small $r$). In order to calculate the central region precisely, we define $x$ by
@@
and $x$ is discretized uniformly with the size $\Delta x$ with $x_{0}$ as the small boundary point. Note that the term $e^{x_{0}}$ is introduced to impose the boundary condition $f_{n\kappa}\left({r=0}\right)=0$ properly, as the following Eq. (30) requiring $f_{n\kappa}\left(r\left(x_{0}\right)\right)=0$. The first-order differential operator for the log mesh reads
@@ IV.3 DNN results
The wave functions of the ground and first five excited states for $\kappa=-1$ are shown in Figs. 3 and 4. In general, both methods can reproduce the exact wave functions, except the peak of the $n=6$ state at large $r$, since the $100\,\mathrm{a.u.}$ box size is too small for the $n=6$ state. The relative errors between DNN wave functions and exact results are shown in Fig. 5. Except around the boundary, the DNN reproduces the exact wave functions. The precisions of wave functions are almost the same level for both methods. Note that due to the existence of nodes, the relative error diverges at those points.
The relative errors of the loss function to the exact energy as a function of epochs for both methods are shown in Fig. 6. As shown in Table 1, the number of epochs to converge for the orthonormal method is one order of magnitude less than the inverse Hamiltonian method, while the time per epoch of the orthonormal method is about twice longer than the inverse Hamiltonian method. The calculation cost of the inverse Hamiltonian method does not depend on $n$; in contrast, the longer calculation for the orthonormal method is required for larger $n$ due to the orthonormalization.
@@ V.2 Discrete representations with uniform mesh
The uniform mesh is adopted to calculate for the Woods-Saxon potential, where $r$ is uniformly distributed with mesh size $\Delta r$. The matrix of the first-order differential operator $\frac{\partial}{\partial r}$ reads
@@ V.3 DNN results
In this section, we use a $20\,\mathrm{fm}$ box with $M=2000$ meshes for the DNN calculation. The general pseudospectral method (GPS) [40, 41] is adopted to calculate benchmark wave functions and energies with the mapping parameter $L=10$ and the number of mesh points $600$.
@@
The $F$- and $G$-components for $\kappa=-1$ of ${}^{208}\mathrm{Pb}$ are shown in Figs. 9 and 10, respectively. Both methods show good agreements with the benchmark for the $F$-component, while the inverse Hamiltonian method gives slightly more accurate results than the orthonormal method for the $G$-component of $n=3$ state. According to Eq. (17), the accuracy of the orthonormal method depends on the accuracy of lower-state wave functions, which means it also depends on the inverse Hamiltonian method which provides the lowest state data.
For the GPS method, which solves the eigenvalue problem of a Dirac Hamiltonian, the computational cost scales as $O\left(M^{3}\right)$, where $M$ is the number of mesh points. For our DNN framework with the inverse Hamiltonian method, the cost is approximately $O\left(XM^{2}\right)$, where $X$ is the number of epochs. As shown in Fig. 6, $X>M$ holds in our example. However, we can use an easier convergence condition to stop the DNN program at a smaller $X$ to reduce the computational cost. Additionally, $X$ remains nearly constant with respect to $M$, which suggests that the DNN method may become more computationally favorable for larger systems requiring larger mesh sizes.
@@
Table 2: Energies of ${}^{16}\mathrm{O}$ below the Fermi energy calculated by the Woods-Saxon potential. The benchmark is obtained by the general pseudospectral method.
@@
$\varepsilon^{\prime}_{n\kappa}$ ($\mathrm{MeV}$) $-45.0$ $-28.0$ $-20.0$
@@
Figure 8: Spectrum of neutrons calculated by the Woods-Saxon potential of ${}^{208}\mathrm{Pb}$. The benchmark is given by the general pseudospectral method and the inverse Hamiltonian method is used for the DNN.
@@ VI Summary
To verify the feasibility of our method, we calculate a hydrogen atom and the Woods-Saxon potentials of ${}^{16}\mathrm{O}$ and ${}^{208}\mathrm{Pb}$. We found that both methods show good agreements with the benchmark with about the same accuracy (up to $0.15\,\%$ difference for the listed states) for the both systems. Additionally, our DNN can consistently reach the desired state regardless of the choice of weight initializers.
One potential extension of our method could involve applications to relativistic many-body systems. For instance, the two-body Dirac equation [42, 43]—which describes relativistic two-particle systems—can, in the center-of-momentum frame, be simplified to a partial differential equation structurally similar to the nonrelativistic Schrödinger equation. In this form, the total energy emerges as an eigenvalue, suggesting that our deep neural network (DNN) approach, with suitable modifications, might be adapted to solve for both ground and excited states in such systems. Additionally, while beyond the scope of the current work, our methodology may hold promise for addressing challenging problems in nuclear physics, such as single-particle resonances in weakly bound nuclei or other systems where traditional computational methods face limitations. These potential applications remain subjects for future investigation.